Print

Print


Public Ticket Saturday 16th February DNS Troubleshooting Parkwest
=================================================================


Ticket Details

 1523402                                                           
 Public Ticket Saturday 16th February DNS Troubleshooting Parkwest 
 open                                                              
 Sat Feb 16 10:00:00 2019                                          
 Sat Feb 16 15:00:00 2019                                          
 YES                                                               

Problem Description

HEAnet require a window to troubleshoot issues that occurred during a planned
migration of hosted services within the Parkwest Datacenter in December 2018.


Problem Effects

During this migration a small number of client services had DNS issues that
interrupted normal behaviour. Following exhaustive review of the day’s events a
root cause for the outage has not been identified so this troubleshooting
window has been agreed, during which it is expected the same client services
will be affected. All clients who will be directly impacted by this
troubleshooting have been informed HEAnet have taken every precaution to
recreate these events in a controlled manner. Rollback plans are in place in
the unlikely event of other, unexpected disruptions to any other client
services.


Last Update Message

Thu Feb 28 11:10:10 2019

Troubleshooting was completed within the window.
Engineers were unable to recreate the core issue encountered in December.
No issues were encountered during the session and the DNS services are now operating on the new network infrastructure as expected.
This session is considered to be successful and further migration plans can now continue.


Affected Clients

n/a


Transaction History

 Date Created              Message                                                                   
 Wed Feb 06 15:31:30 2019  Maintenance START 10:00 Saturday 16th February 2019 Maintenance END -     
                              15:00 Saturday 16th February 2019 HEAnet require a window to              
                              troubleshoot issues that occurred during a planned migration of hosted    
                              services within the Parkwest Datacenter in December 2018. During this     
                              migration a small number of client services had DNS issues that           
                              interrupted normal behaviour. Following exhaustive review of the day’s    
                              events a root cause for the outage has not been identified so this        
                              troubleshooting window has been agreed, during which it is expected       
                              the same client services will be affected. All clients who will be        
                              directly impacted by this troubleshooting have been informed HEAnet       
                              have taken every precaution to recreate these events in a controlled      
                              manner. Rollback plans are in place in the unlikely event of other,       
                              unexpected disruptions to any other client services. Any unflagged        
                              issues during the window should be immediately raised to the HEAnet       
                              NOC on 01 6609040. HEAnet NOC                                             
 Thu Feb 28 11:10:09 2019  Troubleshooting was completed within the window. Engineers were unable    
                              to recreate the core issue encountered in December. No issues were        
                              encountered during the session and the DNS services are now operating     
                              on the new network infrastructure as expected. This session is            
                              considered to be successful and further migration plans can now           
                              continue.